Suppose that N = sqrt x + 4 models the number of cases of an infection, in millions of a disease x years from now. Approx. how many cases of the infection will there be 16 years from now? In approx. how many years will there be 6 million cases?
Answer by stanbon(75887)

Suppose that N = sqrt (x + 4) models the number of cases of an infection, in millions of a disease x years from now.
Approx. how many cases of the infection will there be 16 years from now?
N(16) = sqrt(16+4) = sqrt(20 = 2sqrt(5) is appox. 4.47214 million
----------------------------------------------------------------------
In approx. how many years will there be 6 million cases?
6 = sqrt(x+4)
36 = x+4
x = 32 years
====================
Cheers,
Stan H.